What is the frequency of a microwave in free space whose wavelength is 1.65 cm
Answer #1

Wavelength = 1.65cm = 0.0165 m.

Frequency = speed of light/wavelength = 3*108/0.0165 s-1 = 1.82*1010 s-1.